# MCP-Zero: Proactive Toolchain Construction for LLM Agents from Scratch

### v2

@@ Abstract
Function-calling has enabled large language models (LLMs) to act as tool-using agents, but injecting thousands of tool schemas into the prompt is costly and error-prone. We introduce MCP-Zero, a proactive agent framework that lets the LLM itself decide when and which external tools to retrieve, thereby assembling a task-specific toolchain from scratch. The framework is built upon three components: (1) Proactive Tool Request, where the model emits a structured $\operatorname{tool\_assistant}\right>$ block that explicitly specifies the desired server and task; (2) Hierarchical Vector Routing, a coarse-to-fine retrieval algorithm that first selects candidate servers and then ranks tools within each server based on the semantic similarity; (3) Iterative Proactive Invocation, enabling multi-round, cross-domain toolchain construction with minimal context overhead, and allowing the model to iteratively revise its request when the returned tools are insufficient. To evaluate our approach we also compile MCP-tools, a retrieval dataset comprising 308 MCP servers and 2,797 tools extracted from the official Model-Context-Protocol repository and normalized into a unified JSON schema. Experiments show that MCP-Zero (i) effectively addresses the context overhead problem of existing methods and accurately selects the correct tool from a pool of nearly 3,000 candidates (248.1k tokens); (ii) reduces token consumption by 98% on the APIbank while maintaining high accuracy; and (iii) supports multi-turn tool invocation with consistent accuracy across rounds. The code and dataset will be released soon at https://github.com/xfey/MCP-Zero.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has catalyzed a paradigm shift from pure text understanding and generation to sophisticated tool-using agents capable of interacting with external systems [30, 8, 1]. With the introduction of function-calling mechanisms, LLMs have transcended the boundaries of their parametric knowledge, enabling them to leverage external tools, third-party APIs, and code execution environments to accomplish complex reasoning chains and real-world tasks [21, 26, 14, 18].
However, as the ecosystem of available tools continues to expand exponentially, the conventional approach of injecting comprehensive tool schemas into the system prompt has emerged as a critical bottleneck (Figure LABEL:fig:teasera). Mainstream practices typically embed complete JSON-Schema of all available tools within the system prompt, leading to substantial context overhead, as shown in Figure 1. For instance, the GitHub MCP server encompasses 26 tools requiring over 4,600 tokens, substantially compressing the available context window for actual task content.
To mitigate context window constraints, recent approaches have adopted retrieval-based strategies that match and inject only the most relevant tool based on semantic similarity with user queries [7, 15]. However, these methods face significant limitations when dealing with complex, multi-step tasks that require coordination across multiple domains (Figure LABEL:fig:teaserb). For example, a query like “Debug the file” requires coordinating tools across different domains: $\operatorname{filesystem}$ for file access, $\operatorname{code\_generation}$ for updates, and $\operatorname{command\_execution}$ for debugging. Single-tool retrieval fails to capture such complex workflows, as the initial query cannot determine all required tools across domains.
In summary, while retrieval methods demonstrate efficacy in single-step, single-domain scenarios, they exhibit fundamental limitations in realistic multi-turn, multi-domain agent environments: (1) passive retrieval - external systems select tools based on initial query rather than allowing models to actively express their evolving needs; (2) semantic misalignment between colloquial user inputs and formal API documentation creates distributional mismatches that reduce retrieval precision; and (3) single-round invocation - tool retrieval is performed only once per query, failing to accommodate the progressive refinement of sub-task requirements or iterative correction when initially retrieved tools prove inadequate.
@@
Modern LLMs possess powerful capabilities in chain-of-thought reasoning, self-reflection, and planning. Rather than having external systems passively select tools based on initial queries, we propose letting the model proactively analyze the context, identify capability gaps, and request tools when external assistance is needed. Based on this insight, we introduce MCP-Zero (Figure LABEL:fig:teaserc), a proactive retrieval framework with the following key components:
Proactive Tool Request. Unlike traditional approaches that passively wait for external retrieval, we return the authority of tool requirement specification to the LLM itself. When external tool assistance is needed, the model proactively generates structured tool request declarations in the following format:
@@
This mechanism enables spontaneous need expression while ensuring semantic consistency between requests and tool documentation, avoiding colloquial query issues and improving retrieval quality.
Hierarchical Vector Routing. The system employs two-stage retrieval: first filtering candidate servers by platform requirements, then matching specific tools within selected servers. Only top-k tool descriptions are returned, significantly reducing context overhead.
Iterative Proactive Invocation. The model can initiate multiple tool requests throughout the conversation for different subtasks, enabling cross-server toolchain construction from scratch, while introducing only necessary tools at each step. If returned tools are insufficient, the model can refine its requests and reinitiate retrieval, providing fault tolerance and self-correction capabilities.
To support our approach, we also compile MCP-tools, the first retrieval dataset in the MCP domain, comprising 308 servers and 2,797 tools extracted from the official Model-Context-Protocol repository111github.com/modelcontextprotocol/servers and normalized into a unified JSON schema. Our comprehensive experiments demonstrate that MCP-Zero effectively addresses the context overhead problem while maintaining high accuracy in tool selection from nearly 3,000 candidates, reduces token consumption by 98% compared to existing methods, and supports consistent multi-turn tool invocation across conversation rounds.
Our main contributions are summarized as follows:
  • We propose MCP-Zero, enabling LLMs to proactively analyze tasks, iteratively request tools on demand, and dynamically construct multi-step toolchains with minimal context overhead and high accuracy.
  • We design a hierarchical vector routing mechanism that effectively improves tool retrieval accuracy and context utilization efficiency.
  • We construct and release MCP-tools, a comprehensive retrieval dataset covering 308 servers and 2,797 tools from the official Model-Context-Protocol repository.
  • Extensive experiments demonstrate that our method significantly reduces context overhead while maintaining consistent performance, effectively addressing the context length problem in tool-calling scenarios.
@@ 2.2 Tool Retrieval for LLMs
Figure 2: Illustrative example of MCP-Zero’s iterative proactive invocation. The model progressively identifies capability gaps and requests tools across three domains on-demand.
@@ 3 Method: MCP-Zero
We introduce MCP-Zero through two key aspects: the comprehensive framework design (Section 3.1) and theoretical analysis of proactive retrieval (Section 3.2). Our approach fundamentally shifts from passive tool injection to active tool discovery, enabling LLMs to dynamically construct task-specific toolchains across diverse domains with minimal context overhead.
3.1 Proactive Tool Discovery Framework
MCP-Zero is a proactive agent framework that enables LLMs to dynamically construct task-specific toolchains through on-demand tool retrieval. The framework operates through three core components that work synergistically to address the context overhead and multi-domain coordination challenges of existing approaches.
Overall Workflow. Given a user query such as “Debug my code: src/train.py”, the LLM analyzes the task and autonomously determines when external tool assistance is needed. As shown in Figure 2, the model progressively breaks down the complex debugging task into subtasks: reading the file, analyzing and fixing the code, and validating the fix through execution. At each step, instead of relying on pre-injected tool schemas, the model proactively generates structured tool requests. These requests are processed through hierarchical vector routing to retrieve the most relevant tools, which are then injected into the context for immediate use. The model iteratively repeats this process throughout the conversation, constructing a cross-domain toolchain spanning filesystem operations, code editing, and command execution.
Proactive Tool Request. The foundation of our framework lies in returning tool requirement specification authority to the LLM itself. When the model identifies a capability gap that requires external assistance, it generates a structured request block:
@@ 3.1 Proactive Tool Discovery Framework
Iterative Proactive Invocation. Unlike traditional single-round retrieval approaches, MCP-Zero supports iterative tool discovery throughout task execution. After receiving retrieved tools, the model autonomously evaluates their adequacy for the current subtask. If the returned tools are insufficient or inappropriate, the model can refine its request specification and reinitiate retrieval, providing natural fault tolerance and self-correction capabilities. This iterative process continues until the model determines that either (1) suitable tools have been found and the task can proceed, or (2) no appropriate tools exist and the task should rely on the model’s parametric knowledge.
@@ 3.2 Theoretical Analysis
We provide a theoretical analysis of MCP-Zero’s advantages over traditional approaches through formal probability modeling and complexity analysis.
@@
Problem Formulation. Let $T=\{t_{1},t_{2},...,t_{n}\}$ denote the complete tool collection, $q$ represent the user query, $c$ the conversation context, and $t^{*}$ the optimal tool selection. In traditional approaches, models must select tools from the entire collection simultaneously:
$P_{\text{traditional}}(t^{*}|q,c,T)=\frac{P(q,c|t^{*},T)P(t^{*}|T)}{\sum_{t_{i}\in T}P(q,c|t_{i},T)P(t_{i}|T)}.$ (2)
In contrast, MCP-Zero employs a two-stage process where models first generate structured requests $r$, then retrieve tools based on these requests:
$P_{\text{proposed}}(t^{*}|q,c)=\sum_{r}P(t^{*}|r)P(r|q,c),$ (3)
where $P(r|q,c)$ represents the probability of generating request $r$ given the query and context, and $P(t^{*}|r)$ denotes the probability of retrieving the optimal tool given the request.
Complexity Bottlenecks of Traditional Approaches. Traditional methods face three fundamental limitations that compound as tool collections scale:
Computational Complexity Growth: Models must evaluate all $n$ tools simultaneously, resulting in $O(n)$ computational complexity. This linear growth becomes prohibitive as tool ecosystems expand.
Attention Dilution Effects: Assuming finite attention resources, each tool receives approximately $\frac{1}{n}$ of the model’s attention. Introducing a noise factor $\eta(n)\propto\log(n)$ that increases with collection size, the effective selection probability degrades to:
$P_{\text{noise}}(t^{*}|q,c,T)=P_{\text{traditional}}(t^{*}|q,c,T)\cdot(1-\eta(n)).$ (4)
Semantic Distribution Mismatch: User queries and formal tool descriptions exhibit semantic distribution gaps quantifiable through KL divergence $D_{\text{KL}}(P_{q}\|P_{t})$, which reduces matching accuracy in vector-based retrieval systems.
Advantages of Proactive Request Mechanism. MCP-Zero addresses these limitations through two key mechanisms:
Search Space Compression: The hierarchical routing operates in two stages: first filtering among $m$ servers where $m\ll n$, then matching within the filtered tool subset $T_{\text{filtered}}$ from top-m selected servers. This reduces total complexity from $O(n)$ to $O(m+|T_{\text{filtered}}|)$ where $|T_{\text{filtered}}|\ll n$.
@@
Semantic Consistency Enhancement: Model-generated requests $r$ exhibit higher semantic alignment with tool descriptions compared to raw user queries:
$D_{\text{KL}}(P_{r}\|P_{t})<D_{\text{KL}}(P_{q}\|P_{t}),$ (5)
which enhances vector matching accuracy.
Information-Theoretic Analysis and Probability Gains. From an information-theoretic perspective, traditional methods require decisions over high-entropy spaces. The decision entropy for traditional approaches is:
$H_{\text{traditional}}(T)=-\sum_{t_{i}\in T}P(t_{i})\log P(t_{i})\approx\log(n).$ (6)
MCP-Zero’s hierarchical filtering reduces the effective decision entropy to:
$H_{\text{proposed}}(T_{\text{filtered}})$ $=-\sum_{t_{i}\in T_{\text{filtered}}}P(t_{i})\log P(t_{i})$ (7)
$\approx\log(|T_{\text{filtered}}|)\ll\log(n).$
This entropy reduction translates to measurable probability gains. Defining the probability gain as $\Delta P=P_{\text{proposed}}(t^{*}|q,c)-P_{\text{traditional}}(t^{*}|q,c,T)$. In the traditional uniform, the baseline probability for selecting any specific tool approaches $\frac{1}{n}$. However, the attention dilution effect introduced by processing large tool collections degrades this baseline by the noise factor $\eta(n)$, yielding an effective success probability bounded by $\frac{1}{n}(1-\eta(n))$.
In contrast, MCP-Zero operates over a compressed search space of size $m+|T_{\text{filtered}}|$ where $m\ll n$ and $|T_{\text{filtered}}|\ll n$. More critically, the semantic alignment improvement quantified by the KL divergence reduction introduces a multiplicative enhancement factor $\alpha>1$:
$P_{\text{proposed}}(t^{*}|q,c)$ $\approx\frac{\alpha}{m+|T_{\text{filtered}}|}$ (8)
$P_{\text{traditional}}(t^{*}|q,c,T)$ $\approx\frac{1-\eta(n)}{n}$ (9)
Since $m+|T_{\text{filtered}}|\ll n$ and $\alpha>1$ due to improved semantic consistency, we have $\Delta P>0$ for practical tool collection sizes. Furthermore, as $n$ increases while $m$ and $|T_{\text{filtered}}|$ remain relatively stable (reflecting the hierarchical structure of tool ecosystems), the probability gain $\Delta P$ grows approximately linearly with $n$, demonstrating MCP-Zero’s superior scalability as tool ecosystems expand.
@@ 4 Dataset: MCP-Tools
To support our proactive tool discovery framework and enable comprehensive evaluation, we construct MCP-tools, the first retrieval-oriented dataset in the MCP domain. Unlike existing MCP evaluation frameworks such as MCPBench [13] that focus on server availability and latency testing, our dataset is specifically designed to facilitate semantic tool discovery and retrieval for large language models.
4.1 Data Collection and Scope
We systematically collected MCP server information from the official Model Context Protocol servers repository222github.com/modelcontextprotocol/servers (Tag: 2025.4.28, Commit: ad2d4e6). The repository contains three categories of servers: (1) 20 reference servers maintained by the MCP team, (2) 114 third-party official servers, and (3) 262 community-contributed servers, totaling 396 servers.
Our collection process involved extracting initial metadata including server names, brief descriptions, and repository URLs. Most servers link to GitHub repositories containing comprehensive documentation in README files, which include detailed server descriptions, tool specifications, usage examples, and implementation guidelines.

### v3

# MCP-Zero: Active Tool Discovery for Autonomous LLM Agents

@@ Abstract
Current LLM agents inject thousands of tool schemas into prompts, creating massive context overhead and reducing models to passive tool selectors rather than autonomous agents. We introduce MCP-Zero, an active agent framework that restores tool discovery autonomy to LLMs themselves. Instead of overwhelming models with all available tools, MCP-Zero enables agents to actively identify capability gaps, and request specific tools on-demand, transforming them from large-scale retrievers into genuine autonomous agents. The framework operates through three core mechanisms: (1) Active Tool Request, where models autonomously generate structured requests specifying their exact tool requirements; (2) Hierarchical Semantic Routing, a two-stage algorithm that matches requests to relevant servers and tools through improved semantic alignment; (3) Iterative Capability Extension, enabling agents to progressively build cross-domain toolchains while maintaining minimal context footprint. We also construct MCP-tools, a comprehensive dataset of 308 MCP servers and 2,797 tools from the official Model-Context-Protocol repository. Experiments demonstrate that MCP-Zero preserves agent autonomy while achieving substantial efficiency gains: (i) accurate tool selection from nearly 3k candidates across 248.1k tokens; (ii) 98% reduction in token consumption on APIBank while maintaining high accuracy; and (iii) consistent multi-turn performance that scales with tool ecosystem growth. This work establishes active tool discovery as a fundamental design pattern for scalable autonomous agent systems. The code and dataset is released at https://github.com/xfey/MCP-Zero.
@@ 1 Introduction
Autonomous agents represent a fundamental shift in AI system design: from passive text processors to active decision-makers capable of perceiving their environment, identifying capability gaps, and taking purposeful actions [30, 8, 1]. The emergence of function-calling mechanisms has enabled large language models (LLMs) to transcend their parametric boundaries, leveraging external tools, APIs, and execution environments to accomplish complex real-world tasks [21, 26, 14, 18].
However, current tool integration architectures fundamentally compromise agent autonomy. The dominant paradigm injects comprehensive tool schemas into system prompts, forcing agents into a passive role where they select from pre-defined options rather than actively discovering capabilities as needed (Figure LABEL:fig:teasera). This approach creates two critical problems: massive context overhead and constrained decision autonomy. For instance, the GitHub MCP server requires over 4,600 tokens for 26 tools, while comprehensive tool ecosystems can exceed 248k tokens—effectively transforming capable reasoning models into overwhelmed database query systems.
Recent retrieval-based approaches attempt to reduce context overhead by pre-selecting relevant tools based on semantic similarity with user queries [7, 15]. While these methods address the scaling problem, they perpetuate the fundamental issue of passive tool consumption. Query like “Debug the file” requires filesystem access, code analysis, and command execution (Figure LABEL:fig:teaserb): static retrieval based on initial queries cannot anticipate the evolving tool needs that emerge during task execution. More critically, this paradigm violates a core principle of autonomous agents—the ability to actively shape their environment based on dynamic assessment of their own capabilities.
The limitations of current approaches stem from three architectural constraints that prevent genuine agent autonomy: (1) external decision authority—tool selection is delegated to retrieval systems rather than the agent itself; (2) semantic distribution gaps—user queries and formal tool specifications exist in different semantic spaces, reducing matching precision; and (3) static capability assumptions—tools are selected once rather than discovered iteratively as task understanding evolves.
@@
Toward Active Tool Discovery: True autonomous agents must retain authority over their capability acquisition. Modern LLMs possess sophisticated reasoning, self-reflection, and planning capabilities that enable them to assess their own limitations and articulate specific tool requirements. Rather than constraining agents to pre-selected tool sets, we propose active tool discovery—a paradigm where agents autonomously identify capability gaps and request appropriate tools on-demand. Based on this principle, we introduce MCP-Zero (Figure LABEL:fig:teaserc), an active agent framework that restores tool discovery autonomy to LLMs through three core mechanisms:
Active Tool Request. Instead of passive tool consumption, agents generate structured requests that specify their exact requirements:
@@
This approach ensures semantic alignment between agent needs and tool documentation while preserving decision autonomy.
Hierarchical Semantic Routing. A two-stage matching algorithm first filters candidate servers by platform requirements, then ranks tools within selected servers based on semantic similarity. This hierarchical approach reduces search complexity while maintaining precision.
Iterative Capability Extension. Agents can discover and integrate tools throughout task execution, building cross-domain capabilities dynamically. When initial tools prove insufficient, agents can refine their requests and discover alternatives, providing natural fault tolerance.
To support systematic evaluation, we also construct MCP-tools, a comprehensive dataset comprising 308 servers and 2,797 tools from the official Model-Context-Protocol repository111github.com/modelcontextprotocol/servers. Experiments demonstrate that MCP-Zero maintains agent autonomy while achieving substantial efficiency gains: 98% reduction in token consumption with preserved accuracy across multi-turn conversations and large-scale tool ecosystems.
Our main contributions establish active tool discovery as a fundamental design pattern for autonomous agent systems:
  • We propose MCP-Zero, enabling agents to maintain decision autonomy through active tool discovery, transforming them from passive selectors to autonomous capability architects.
  • We design hierarchical semantic routing that preserves semantic alignment between agent requests and tool specifications while reducing computational complexity.
  • We construct and release MCP-tools, providing the research community with a comprehensive evaluation framework for tool discovery systems.
  • We demonstrate that active discovery paradigms scale effectively with tool ecosystem growth while maintaining the autonomous decision-making that defines genuine agent systems.
@@ 2.2 Tool Retrieval for LLMs
Figure 2: Illustrative example of MCP-Zero’s iterative active invocation. The model progressively identifies capability gaps and requests tools across three domains on-demand.
@@ 3 Method: MCP-Zero
We introduce MCP-Zero through two key aspects: the comprehensive framework design (Section 3.1) and theoretical analysis of active retrieval (Section 3.2). Our approach fundamentally shifts from passive tool injection to active tool discovery, enabling LLMs to dynamically construct task-specific toolchains across diverse domains with minimal context overhead.
3.1 Active Tool Discovery Framework
MCP-Zero is a active agent framework that enables LLMs to dynamically construct task-specific toolchains through on-demand tool retrieval. The framework operates through three core components that work synergistically to address the context overhead and multi-domain coordination challenges of existing approaches.
Overall Workflow. Given a user query such as “Debug my code: src/train.py”, the LLM analyzes the task and autonomously determines when external tool assistance is needed. As shown in Figure 2, the model progressively breaks down the complex debugging task into subtasks: reading the file, analyzing and fixing the code, and validating the fix through execution. At each step, instead of relying on pre-injected tool schemas, the model actively generates structured tool requests. These requests are processed through hierarchical vector routing to retrieve the most relevant tools, which are then injected into the context for immediate use. The model iteratively repeats this process throughout the conversation, constructing a cross-domain toolchain spanning filesystem operations, code editing, and command execution.
Active Tool Request. The foundation of our framework lies in returning tool requirement specification authority to the LLM itself. When the model identifies a capability gap that requires external assistance, it generates a structured request block:
@@ 3.1 Active Tool Discovery Framework
Iterative Active Invocation. Unlike traditional single-round retrieval approaches, MCP-Zero supports iterative tool discovery throughout task execution. After receiving retrieved tools, the model autonomously evaluates their adequacy for the current subtask. If the returned tools are insufficient or inappropriate, the model can refine its request specification and reinitiate retrieval, providing natural fault tolerance and self-correction capabilities. This iterative process continues until the model determines that either (1) suitable tools have been found and the task can proceed, or (2) no appropriate tools exist and the task should rely on the model’s parametric knowledge.
@@ 3.2 Theoretical Analysis
We provide a theoretical analysis of MCP-Zero’s advantages over traditional approaches through formal modeling that connects to established theories in active learning and information acquisition.
@@
Problem Formulation. Let $T=\{t_{1},t_{2},...,t_{n}\}$ denote the complete tool collection, $q$ represent the user query, $s_{t}$ the current conversation state, and $t^{*}$ the optimal tool selection. Traditional approaches require simultaneous evaluation over the entire collection:
$P_{\text{passive}}(t^{*}|q,T)=\frac{P(q|t^{*},T)P(t^{*}|T)}{\sum_{t_{i}\in T}P(q|t_{i},T)P(t_{i}|T)}$ (2)
MCP-Zero employs active information acquisition where agents generate requests $r$ based on their current state and capability assessment:
$P_{\text{active}}(t^{*}|s_{t})=\sum_{r}P(t^{*}|r)P(r|s_{t})$ (3)
where $P(r|s_{t})$ represents the agent’s ability to articulate its needs given current understanding.
Active Information Acquisition. Tool request generation can be modeled as an active learning process where agents select actions to maximize information gain about task completion:
$r^{*}$ $=\arg\max_{r}I(T^{*};r|s_{t})$ (4)
$=\arg\max_{r}[H(T^{*}|s_{t})-H(T^{*}|r,s_{t})]$
where $I(T^{*};r|s_{t})$ represents the mutual information between the optimal tool set and the request. This formulation captures the essence of active discovery: agents actively reduce uncertainty about their tool requirements.
Scalability Analysis. Traditional methods face fundamental scalability barriers:
Search Space Complexity: Passive approaches must process all $n$ tools with $O(n)$ complexity. Active approaches first filter among $m$ servers ($m\ll n$), then match within filtered subsets, reducing complexity to $O(m+k)$ where $k$ is the average tools per selected server.
Attention Distribution: With finite cognitive resources, passive methods distribute attention as $\frac{1}{n}$ per tool, degraded by noise factor $\eta(n)\propto\log(n)$. Active methods concentrate attention on relevant subsets, maintaining effectiveness as $\frac{1}{k}$ where $k\ll n$.
@@
Semantic Alignment Advantage. The core advantage of active requests lies in improved semantic consistency. Agent-generated requests $r$ exhibit stronger alignment with tool documentation compared to raw user queries $q$:
$\text{Alignment}(r,t)=\cos(\mathbf{e}_{r},\mathbf{e}_{t})>\cos(\mathbf{e}_{q},\mathbf{e}_{t})$ (5)
where $\mathbf{e}_{r}$, $\mathbf{e}_{q}$, $\mathbf{e}_{t}$ represent embeddings of request, query, and tool description respectively. This improvement stems from agents operating in the same semantic space as tool documentation.
Iterative Information Gain. Unlike single-shot retrieval, active discovery enables cumulative information acquisition over $k$ iterations:
$I_{\text{total}}=\sum_{i=1}^{k}I(T^{*};r_{i}|s_{i-1})-\lambda\cdot\text{Cost}(r_{i})$ (6)
where $\lambda$ represents the context overhead per request. This formulation captures the trade-off between information gain and computational efficiency that MCP-Zero optimizes.
In summary, the active paradigm provides measurable advantages:
  • Complexity Reduction: From $O(n)$ to $O(m+k)$ where $m+k\ll n$.
  • Semantic Consistency: Direct embedding alignment in tool description space.
  • Information Efficiency: Targeted uncertainty reduction rather than exhaustive search.
  • Adaptive Capability: State-dependent tool discovery that evolves with task understanding.
These theoretical foundations explain the empirical performance gains observed in our experiments: 98% token reduction with maintained accuracy reflects the efficiency of targeted information acquisition over exhaustive tool enumeration.
@@ 4 Dataset: MCP-Tools
To support our active tool discovery framework and enable comprehensive evaluation, we construct MCP-tools, the first retrieval-oriented dataset in the MCP domain. Unlike existing MCP evaluation frameworks such as MCPBench [13] that focus on server availability and latency testing, our dataset is specifically designed to facilitate semantic tool discovery and retrieval for large language models.
4.1 Dataset Construction
Our dataset is constructed with the following steps:
Data Collection. We systematically collected MCP server information from the official Model Context Protocol repository222github.com/modelcontextprotocol/servers (Tag: 2025.4.28, Commit: ad2d4e6), encompassing 396 servers across three categories: 20 reference implementations, 114 third-party official servers, and 262 community contributions. This collection represents the current state of the MCP ecosystem and its diversity of tool capabilities.
Quality Assurance and Filtering. We implemented rigorous filtering to ensure dataset completeness and semantic richness. Servers were retained only if they provided: (1) MCP-compliant tool definitions, (2) comprehensive documentation enabling semantic understanding, and (3) sufficient detail for meaningful retrieval evaluation. This process yielded 308 high-quality servers representing the core of the MCP ecosystem.